# Cost-aware Portfolios in a Large Universe of Assets

### v1

@@ Abstract
This paper considers the finite horizon portfolio rebalancing problem in terms of mean-variance optimization, where decisions are made based on current information on asset returns and transaction costs. The study’s novelty is that the transaction costs are integrated within the optimization problem in a high-dimensional portfolio setting where the number of assets is larger than the sample size. We propose portfolio construction and rebalancing models with nonconvex penalty considering two types of transaction cost, the proportional transaction cost and the quadratic transaction cost. We establish the desired theoretical properties under mild regularity conditions. Monte Carlo simulations and empirical studies using S&P 500 and Russell 2000 stocks show the satisfactory performance of the proposed portfolio and highlight the importance of involving the transaction costs when rebalancing a portfolio.
Keywords: High-dimensional Portfolio Optimization, Optimal Rebalancing, SCAD, Transaction Costs.
@@ 1 Introduction
where $\bm{w}$ is the weight vector of the assets in the portfolio, $\gamma$ is the inverse of the risk aversion parameter, $\bm{\mu}=(\mu_{1},\cdots,\mu_{p})^{\top}$ is the mean vector of the excess returns of $p$ assets and $\bm{\Sigma}$ is the covariance matrix of the excess returns. Here we assume that $\sum_{i=1}^{p}w_{i}=1$, so the portfolio is fully-invested on the pool of $p$ risky assets. Samuelson (1975) extends the Markowitz (1952) model, which considers discrete multi-periods portfolio allocation. The above method has an elegant analytic solution that depends only on the excess returns’ mean and covariance matrix. The solution to (1.1) is $\bm{w}^{*}=c_{1}\bm{\Sigma}^{-1}\bm{\mu}-c_{2}\bm{\Sigma}^{-1}\mathbf{1}$, where $c_{1}$ and $c_{2}$ depending only on $\bm{\Sigma}$ and $\bm{\mu}$. However, in modern portfolio management, portfolios commonly exhibit high-dimensional characteristics, introducing new challenges to the traditional framework. Notably, the inverse of the sample covariance matrix is not well defined when the number of assets is larger than the sample size.
Besides the difficulties of covariance matrix estimation in high-dimensional portfolio allocation, transaction costs also make the classical Markowitz mean-variance model sub-optimal. In most empirical studies, transaction costs are retrospectively incorporated by analyzing a portfolio strategy’s performance under given-sized transaction costs. However, in financial practice, the costs of the portfolio are considered during the decision-making process, meaning the costs are a part of the optimization problem. For minor trades that have negligible effects on market prices, transaction cost is generally assumed to be proportional to the trading amount. For larger trades, the literature usually assumes that they have a nonlinear impact on the market price, resulting in quadratic transaction costs (Olivares-Nadal and DeMiguel, 2018). Hautsch and Voigt (2019) illustrated that when incorporating transaction costs into the optimization problem, quadratic transaction costs can be understood as causing the covariance matrix to contract towards a diagonal matrix and a mean shift proportionate to current holdings. Proportional transaction cost, on the other hand, is proportional to the sum of absolute rebalancing, which imposes a stronger penalization on turnover and thus reduces the volume and frequency of rebalancing. Some studies have provided empirical evidence that strategies incorporating transaction fees consistently demonstrate superior performance compared to the minimum variance benchmark methods (Du et al., 2022; Hautsch and Voigt, 2019).
@@ 1.1 Main Takeaways
This paper considers the problem faced by a real mutual fund or sophisticated investor. Transaction cost is an integral part of portfolio management, which is passed on to investors who can get returns on the net of such costs. We propose an innovative method for constrained portfolio selection tailored for high-dimensional portfolios, taking into account two different types of transaction costs. Moreover, our approach allows for adaptive weight adjustments in multi-stage scenarios. In other words, we construct the optimal portfolio that considers cost of rebalancing in the optimization procedure.
The main contributions of our work are threefold. Firstly, we propose cost-aware portfolio estimators with nonconvex penalty for portfolio construction and rebalancing, and two types of transaction costs are considered in the proposed estimators. The nonconvex penalization can improve the estimation accuracy in contrast with the $\ell_{1}$ penalization method. We show the theory of the proposed portfolio in the construction and rebalancing stages under both cost functions. Secondly, we adopt the local linear approximation (LLA) algorithm (Zou and Li, 2008) to solve the nonconvex optimization problem and develop the theoretical foundations of the proposed portfolio with LLA algorithm. Thirdly, we introduce data-dependent penalization models for large portfolios in both construction and rebalancing periods that consider transaction costs and demonstrate its superior performance in simulation studies and real data examples.
@@ 1.2 Brief Literature Review
Our paper connects to the literature on covariance learning and high-dimensional portfolio optimization. Fan et al. (2015) studied a factor-based risk estimator under many assets and introduced a high-confidence level upper bound to assess the estimation. Lam and Feng (2018) proposed a nonparametrically eigenvalue-regularized integrated covariance matrix estimator. Fan et al. (2019) constructed a robust covariance estimation under the approximate factor model with observed factors. Hafner et al. (2020) developed a Kronecker product model for the covariance matrix estimation and established central limit theorems for such estimator. Ren et al. (2015), Fan and Lv (2016) and Ren et al. (2019) conducted studies on the estimation and statistical inference of sparse precision matrix under the assumption of conditional normal distribution. Nevertheless, practical challenges arise in verifying the normality of high-dimensional data. Wang et al. (2020) introduced a refitted cross-validation estimation method for a high-dimensional precision matrix utilizing its Cholesky decomposition. Other recent works include but are not limited to Engle et al. (2019); Cai et al. (2020); So et al. (2022); Ding et al. (2021). The estimates in these works above converge to the true covariance matrix or precision matrix under certain regularity conditions. However, whether the derived portfolio allocation provides the optimal strategy with cost consideration is still unclear.
@@
Another strand of literature considers various constraints on large portfolio optimization. Jagannathan and Ma (2003) proposed no-short-selling constraints to reduce the portfolio risks. Inspired by Jagannathan and Ma (2003), DeMiguel et al. (2009) imposed the norm constraints to the portfolios. Fan et al. (2012) proposed a gross-exposure constraint method to modify the Markowitz mean-variance optimization problem in a single period to reduce the sensitivity of the resulting allocation to input vectors. Du et al. (2022) showed that under mild conditions, solving the $\ell_{1}$ constrained problem yields the same expected utility as solving the $\ell_{0}$ constrained problem, enabling effective implementation of cardinality constraints through $\ell_{1}$ constraints. Zou (2006); Zhao and Yu (2006); Gai et al. (2013) suggested that $\ell_{1}$ regularization methods demonstrate support recovery property under restrictive regularity conditions while being prone to bias for large coefficients. In contrast, nonconvex penalization methods such as SCAD (Fan and Li, 2001) and MCP (Zhang, 2010) exhibit favorable theoretical properties without relying on restrictive conditions and effectively mitigate the estimation bias associated with Lasso.
@@
On the other hand, the literature on transaction costs and portfolio optimization is way more scarce, especially when large dimensions are considered. One exception is the important work by Ledoit and Wolf (2024), where the authors also consider transaction costs at the optimization stage when constructing mean-variance portfolios. However, their paper focuses only on the empirical part and provides no theoretical guidance. Furthermore, the authors do not consider extra regularization terms to discipline the portfolio in large dimensions. In our opinion, our paper complements the analysis in Ledoit and Wolf (2024).
@@ 1.3 Organization of the Paper
The remaining of the article is organized as follows. Section 2 proposes the constrained nonconvex penalization model for high-dimensional portfolios in both construction and rebalacing stages with consideration of transaction cost. We also introduce the LLA to solve the proposed model. In Section 3, we conduct a comprehensive numerical study to examine the performance of the proposed method. Empirical analysis is included in Section 4. We summarize our contribution in Section 5.
Notations: We use the standard notational convention that constants $c_{1}$,$c_{2}$, etc., represent universal positive constants, with values that may change from line to line. We write $f(n)=O(g(n))$ to mean $f(n)\leq cg(n)$ for some constant $c\in(0,\infty)$. For a matrix $\bm{M}$, we write $|||\bm{M}|||_{\infty}$ to denote the $L_{\infty}$ norm, $\|\bm{M}\|_{\max}=$ max$|m_{ij}|$ to denote its entry-wise $L_{\infty}$ norm, and $|||\bm{M}|||_{2}$ to denote the spectral norm. $\lambda_{\min}(\bm{M})$ and $\lambda_{\max}(\bm{M})$ are minimum and maximum eigenvalues of $\bm{M}$. For a vector $\bm{v}=(v_{j})_{j=1}^{p}\in\mathbb{R}^{p}$, we write $\|\bm{v}\|_{k}$, $k=1,2$ to denote its $L_{k}$ norm, $\|\bm{v}\|_{0}$ is the number of nonzero entries of $\bm{v}$, and $L_{\infty}$ norm is $\|\bm{v}\|_{\infty}=$max$|v_{j}|$. For subsets $\mathcal{A}\subset[1,\ldots,p]$ and $\mathcal{B}\subset[1,\ldots,p]$, let $\bm{M}_{\mathcal{A},\mathcal{B}}$ denote the submatrix consisting of the rows indexed by $\mathcal{A}$ and the columns indexed by $\mathcal{B}$, and $\bm{v}_{\mathcal{A}}$ denote the subvector containing elements of $\bm{v}$ indexed by the set $\mathcal{A}$. The cardinality of $\mathcal{A}$ is represented as $|\mathcal{A}|$. The symbol $\odot$ denotes the element-by-element multiplication of two vectors. For a function $h:\mathbb{R}^{p}\rightarrow\mathbb{R}$, we write $\nabla h$ to denote a gradient or subgradient.
@@ 2.1 Cost-aware Portfolio Estimator
We consider an investor who wants to allocate the wealth among $p$ risky assets and assume the asset returns to be stationary. We denote the expected asset returns as
$\bm{\mu}=\left(\mu_{1},\mu_{2},\ldots,\mu_{p}\right)^{\top}.$
Suppose an investor enters the market at day 1 with an initial wealth, allocates the wealth (constructing the first portfolio) at day $n+1$, where $n\in\mathcal{N}^{+}$ is a positive number of trading days between two consecutive rebalances, and reallocates the wealth  [^1] For simplicity but without loss of generality, we assume the investor rebalances the portfolio at same frequency, e.g., monthly, quarterly, annually. at the beginning of each of the following $m-1$ consecutive periods, where $m$ is the total number of portfolio formation decisions. At every decision day $d$, $d=n+1,2n+1,\ldots,mn+1$, we use the historical returns $[\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}]^{\top}\in\mathbb{R}^{n\times p}$, that contains observed returns of asset $j$ on its $j$-th column to estimate assets return and risk. Figure 1 illustrates the decision process.
@@
and the weights are taken at the beginning of period $t$. Considering the transaction cost under the Markowitz mean-variance framework (1.1), the optimal cost-aware portfolio at first construction stage is defined as
$\bm{w}^{*}_{1}:=\mathop{\arg\min}_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\bm{\Sigma}_{1}\bm{w}-\gamma\bm{w}^{\top}\bm{\mu}_{1}+\mathbf{C}_{1}(\bm{w}),$ (2.1)
where $\mathbf{C}_{1}(\bm{w})$ is the transaction cost, depending on the trading activity (changes in portfolio weights). Similarly, the optimal cost-aware portfolio at reallocation stage is defined as
$\bm{\delta}^{*}_{t}:=\mathop{\arg\min}_{\bm{\delta}\in\mathbf{R}^{p},\bm{\delta}^{\top}\mathbf{1}=0}\bm{\delta}^{\top}\bm{\Sigma}_{t}\bm{\delta}-\gamma\bm{\delta}^{\top}\bm{\mu}_{t}+\mathbf{C}_{t}(\bm{\delta}),$ (2.2)
where the weight difference vector $\bm{\delta}_{t}=\bm{w}^{*}_{t}-\bm{w}^{+}_{t-1}$ and $\bm{w}^{+}_{t-1}$ means the pre-rebalancing weight as the value of portfolio changes over the holding period, and $\bm{w}^{+}_{t-1}=(f_{n}\circ f_{n-1}\circ\dots\circ f_{1})(\bm{w}_{t-1})$ for $t\geq 2$, where $f_{i}(\bm{w}):=\frac{\bm{w}\odot(\mathbf{1}+\bm{R}_{(t-2)n+i})}{1+\bm{w}^{\top}\bm{R}_{(t-2)n+i}}$. This method is thus called the cost-aware portfolio estimator (CAPE). This paper proposes a new estimator targeting $\bm{w}^{*}_{t}$ under the high-dimensional model setting.

### v2

@@ Abstract
This paper proposes a finite-horizon (multi-period) mean-variance portfolio estimator, in which rebalancing decisions are based on current information about asset returns and transaction costs. The novelty of this study stems from integrating transaction costs into the decision process within a high-dimensional portfolio setting, where the number of assets exceeds the sample size. We define the optimal cost-aware portfolio and propose novel models for its construction and rebalancing. Our approach incorporates a nonconvex penalty and explicitly accounts for both proportional and quadratic transaction costs. We establish that the estimators derived from the proposed construction and rebalancing models, as well as the corresponding in-sample and out-of-sample Sharpe ratio estimators, consistently converge to those of the optimal cost-aware portfolio. Monte Carlo simulations and empirical studies using S&P 500 and Russell 2000 stocks show the satisfactory performance of the proposed portfolio and highlight the importance of incorporating transaction costs during rebalancing.
Keywords: High-dimensional Portfolio Optimization, Mean-variance Model, Optimal Rebalancing, Transaction Costs.
@@ 1 Introduction
where $\bm{w}$ is the weight vector of the assets in the portfolio, $\gamma$ is the inverse of the risk aversion parameter, $\bm{\mu}=(\mu_{1},\cdots,\mu_{p})^{\top}$ is the mean vector of the excess returns of $p$ assets and $\bm{\Sigma}$ is the covariance matrix of the excess returns. Here we assume that $\sum_{i=1}^{p}w_{i}=1$, so the portfolio is fully-invested on the pool of $p$ risky assets. Samuelson (1975) extends the Markowitz (1952)model to consider discrete multi-period portfolio allocation. This approach has an elegant analytical solution that depends only on the mean and covariance matrix of the excess returns. The solution to (1.1) is $\bm{w}^{*}=c_{1}\bm{\Sigma}^{-1}\bm{\mu}-c_{2}\bm{\Sigma}^{-1}\mathbf{1}$, where $c_{1}$ and $c_{2}$ depend only on $\bm{\Sigma}$ and $\bm{\mu}$. However, in modern portfolio management, portfolios commonly exhibit high-dimensional characteristics, introducing new challenges to the traditional framework. Notably, the inverse of the sample covariance matrix is not well defined when the number of assets is larger than the sample size.
Besides the difficulties of covariance matrix estimation in high-dimensional portfolio allocation, transaction costs also make the classical Markowitz mean-variance model sub-optimal. In most empirical studies, transaction costs are retrospectively incorporated by analyzing a portfolio strategy’s performance under given-sized transaction costs. However, in financial practice, the costs of the portfolio are considered during the decision-making process, meaning the costs are an integral part of the rebalancing. Transaction costs are generally assumed to be proportional to the trading amount for minor trades with negligible market impact. For larger trades, the literature usually assumes that they have a nonlinear impact on the market price, resulting in quadratic transaction costs (Olivares-Nadal and DeMiguel, 2018). Hautsch and Voigt (2019) illustrated that when incorporating transaction costs into the optimization problem, quadratic transaction costs can be understood as causing the covariance matrix to contract towards a diagonal matrix and a mean shift proportionate to current holdings. Proportional transaction costs scale with the sum of the absolute values of rebalancing amounts, penalizing turnover more heavily and thereby reducing rebalancing volume and frequency. Some studies have provided empirical evidence that strategies incorporating transaction fees consistently demonstrate superior performance compared to the minimum variance benchmark methods (Du et al., 2022; Hautsch and Voigt, 2019).
@@ 1.1 Main Takeaways
This paper considers the problem faced by a real mutual fund or sophisticated investor. Transaction costs are an integral part of portfolio management and are passed on to investors, who receive returns net of such costs. We propose an innovative method for constrained portfolio selection tailored for high-dimensional portfolios, taking into account two different types of transaction costs. Moreover, our approach allows for adaptive weight adjustments in multi-stage scenarios. Specifically, at each decision point, the portfolio weights are rebalanced in a data-driven manner based on the current holdings and the return information from the previous stage. In other words, we construct the optimal portfolio that considers the cost of rebalancing, which provides guidance to cost-aware investors.
The main contributions of our work are threefold. Firstly, we propose cost-aware optimal portfolio estimators for new construction and subsequent rebalancing in a high-dimensional framework. To our knowledge, such a portfolio estimator in high dimensions appears to be novel. Also, the nonconvex penalization employed in our approach can improve the estimation accuracy compared to the $\ell_{1}$ penalization method. We establish theoretical properties for the proposed portfolio in both the construction and rebalancing stages under each cost function. We show that, under both cost functions, the estimators obtained in the construction and rebalancing stages—as well as the corresponding in-sample and out-of-sample Sharpe ratio estimators—consistently converge to those of the optimal cost-aware portfolio. Secondly, we adopt the local linear approximation (LLA) algorithm (Zou and Li, 2008) to solve the nonconvex optimization problem and develop the theoretical foundations of the proposed portfolio with LLA algorithm. Thirdly, we introduce data-dependent penalization for investors facing a large universe of assets in each period while considering transaction costs. We demonstrate the superior performance of the proposed portfolio through extensive simulation studies and empirical evaluations. In particular, we compare it against a wide range of existing portfolio construction methods. For the real data analysis, we use stock return data from the S&P 500 and Russell 2000 indices over the period 2017–2020.
@@ 1.2 Brief Literature Review
Our paper connects to the literature on covariance learning and high-dimensional portfolio optimization.
@@
Fan et al. (2015) studied a factor-based risk estimator under many assets and introduced a high-confidence level upper bound to assess the estimation. Lam and Feng (2018) proposed a nonparametrically eigenvalue-regularized integrated covariance matrix estimator. Fan et al. (2019) constructed a robust covariance estimation under the approximate factor model with observed factors. Hafner et al. (2020) developed a Kronecker product model for the covariance matrix estimation and established central limit theorems for such estimator. Ren et al. (2015), Fan and Lv (2016) and Ren et al. (2019) conducted studies on the estimation and statistical inference of sparse precision matrix under the assumption of conditional normal distribution. Nevertheless, practical challenges arise in verifying the normality of high-dimensional data. Wang et al. (2020) introduced a refitted cross-validation estimation method for a high-dimensional precision matrix utilizing its Cholesky decomposition. Other recent works include but are not limited to Engle et al. (2019); Cai et al. (2020); So et al. (2022); Ding et al. (2021).
To enhance out-of-sample performance, Luo et al. (2024) proposed an iterative optimization framework based on almost second-degree stochastic dominance (ASSD), which improves convergence efficiency and empirically yields higher out-of-sample returns compared to traditional dominance and mean-variance approaches.
The estimators in the works above converge to the true covariance matrix or precision matrix under certain regularity conditions. However, whether the derived portfolio allocation provides the optimal strategy with cost consideration is still unclear.
Another strand of literature considers various constraints on large portfolio optimization. Jagannathan and Ma (2003) proposed no-short-selling constraints to reduce the portfolio risks. Inspired by Jagannathan and Ma (2003), DeMiguel et al. (2009) imposed the norm constraints to the portfolios. In particular, they developed a unified framework for portfolio optimization by constraining the norm of the portfolio weights, which effectively mitigates estimation error and encompasses several existing strategies such as shrinkage-based estimators and the 1/N portfolio.
@@
Fan et al. (2012) proposed a gross-exposure constraint method to modify the Markowitz mean-variance optimization problem in a single period to reduce the sensitivity of the resulting allocation to input vectors. Du et al. (2022) showed that under mild conditions, solving the $\ell_{1}$ constrained problem yields the same expected utility as solving the $\ell_{0}$ constrained problem, enabling effective implementation of cardinality constraints through $\ell_{1}$ constraints. Zou (2006); Zhao and Yu (2006); Gai et al. (2013) suggested that $\ell_{1}$ regularization methods demonstrate the support recovery property under restrictive regularity conditions while being prone to bias for large coefficients. In contrast, nonconvex penalization methods such as SCAD (Fan and Li, 2001) and MCP (Zhang, 2010) exhibit favorable theoretical properties without relying on restrictive conditions and effectively mitigate the estimation bias associated with Lasso.
In addition, Ban et al. (2018) proposed performance-based regularization and cross-validation techniques for portfolio optimization, casting the problem into a robust optimization framework and showing that their method outperforms traditional and regularized approaches across several Fama-French datasets.
@@
On the other hand, the literature on transaction costs and portfolio optimization in large dimensions is significantly less abundant. Most of the papers in Mathematical Finance and Operations Research focus on low-dimensional settings. See, for example, Lobo et al. (2007) and Olivares-Nadal and DeMiguel (2018). One exception is the important work by Ledoit and Wolf (2025), where the authors also consider transaction costs at the optimization stage when constructing mean-variance portfolios. However, their paper focuses only on the empirical part and provides no theoretical guidance. Furthermore, the authors do not consider extra regularization terms to discipline the portfolio in large dimensions. In our opinion, our paper complements the analysis in Ledoit and Wolf (2025).
From an algorithmic perspective, Best and Hlouskova (2005) propose an efficient method for portfolio optimization with transaction costs, which reduces a high-dimensional constrained problem into a sequence of lower-dimensional subproblems by implicitly incorporating transaction cost effects. Similarly, Perold (1984) provides a computationally efficient algorithm for large-scale mean-variance portfolio optimization with transaction costs and factor-based models, enabling parametric solutions under practical resource constraints.
In the dynamic setting, Brown and Smith (2011) study multi-period portfolio optimization with transaction costs and propose heuristic strategies that perform close to optimal, supported by dual-based upper bounds derived from information relaxation.
@@ 1.3 Organization of the Paper
The remainder of the article is organized as follows. Section 2 proposes the constrained nonconvex penalization model for high-dimensional portfolios in both construction and rebalacing stages with consideration of transaction cost. We also introduce the LLA to solve the proposed model. In Section 3, we conduct a comprehensive numerical study to examine the performance of the proposed method. Empirical analysis is included in Section 4. We summarize our contribution in Section 5.
Notations: We use the standard convention that constants $c_{1}$, $c_{2}$, etc., denote universal positive constants whose values may vary from line to line. We write $f(n)=O(g(n))$ if $f(n)\leq cg(n)$ for some constant $c>0$. For a matrix $\bm{M}$, we let $|\bm{M}|||_{\infty}$ denote the $L_{\infty}$ operator norm, $\|\bm{M}\|_{\max}=\max|m_{ij}|$ denote the entrywise $L_{\infty}$ norm, and $|\bm{M}|||_{2}$ denote the spectral norm. The minimal and maximal eigenvalues of $\bm{M}$ are denoted by $\lambda_{\min}(\bm{M})$ and $\lambda_{\max}(\bm{M})$, respectively. For a vector $\bm{v}=(v_{j})_{j=1}^{p}\in\mathbb{R}^{p}$, we use $\|\bm{v}\|_{k}$ ($k=1,2$) for its $L_{k}$ norm, $\|\bm{v}\|_{0}$ for the number of nonzero entries, and $\|\bm{v}\|_{\infty}=\max|v_{j}|$ for the $L_{\infty}$ norm.
Given index sets $\mathcal{A},\mathcal{B}\subset\{1,\ldots,p\}$, we write $\bm{M}_{\mathcal{A},\mathcal{B}}$ for the submatrix of $\bm{M}$ with rows indexed by $\mathcal{A}$ and columns by $\mathcal{B}$, and $\bm{v}_{\mathcal{A}}$ for the subvector of $\bm{v}$ with entries in $\mathcal{A}$. The cardinality of $\mathcal{A}$ is denoted $|\mathcal{A}|$. The symbol $\odot$ represents element-wise multiplication. For a function $\mathbb{R}^{p}\rightarrow\mathbb{R}, h\nabla h$ denotes its gradient or subgradient. To simplify notation, when referring to time-specific quantities such as $\bm{\mu}_{t}$ and $\bm{\Sigma}_{t}$, we write $\bm{\mu}_{\mathcal{A}_{t}}$ and $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$ in place of the more precise $\bm{\mu}_{t,\mathcal{A}_{t}}$ and $\bm{\Sigma}_{t,\mathcal{A}_{t},\mathcal{A}_{t}}$ whenever the context is clear.
@@ 2 Methodology
2.1 Preliminary
To contextualize our proposed approach, we review three foundational models in the mean-variance portfolio optimization literature. These models serve as natural precursors to our method and are also employed as benchmark comparisons in our simulation studies.
  • 1.
    The classical Markowitz mean-variance model. Originally introduced by Markowitz (1952), this model aims to balance risk and return by minimizing portfolio variance for a given level of expected return. It is formulated as:
    $\min_{\bm{w}}\ \bm{w}^{\top}\widehat{\bm{\Sigma}}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}},\quad\text{subject to }\sum_{i=1}^{p}w_{i}=1,$ (2.1)
    where $w_{i}$ denotes the $i$th element of the portfolio weight vector $\bm{w}$, $\widehat{\bm{\Sigma}}$ is the estimated covariance matrix, and $\widehat{\bm{\mu}}$ is the vector of estimated expected returns. Despite its theoretical elegance, the Markowitz model often suffers from poor empirical performance in high-dimensional settings due to estimation errors in the covariance matrix.
  • 2.
    The penalized mean-variance model with gross-exposure constraints. Fan et al. (2012) propose a regularized version of the classical mean-variance framework by incorporating an $\ell_{1}$ penalty on the portfolio weights to control gross exposure and improve stability in high-dimensional settings. The optimization problem is:
    $\min_{\bm{w}}\ \bm{w}^{\top}\widehat{\bm{\Sigma}}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}+\lambda\|\bm{w}\|_{1},\quad\text{subject to }\sum_{i=1}^{p}w_{i}=1,$ (2.2)
    where $\lambda>0$ is a tuning parameter that regulates the portfolio’s gross exposure. This formulation mitigates the accumulation of estimation errors in large portfolios and can outperform no-short-sale constraints by allowing limited short positions.
  • 3.
    The mean-variance model with transaction costs. Hautsch and Voigt (2019) extend the mean-variance paradigm to account for transaction costs explicitly. By penalizing turnover directly, their model promotes more stable allocations while implicitly regularizing the covariance matrix. The corresponding optimization problem is:
    $\min_{\bm{w}}\ \bm{w}^{\top}\widehat{\bm{\Sigma}}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}+\mathbf{C}(\bm{w}),\quad\text{subject to }\sum_{i=1}^{p}w_{i}=1,$ (2.3)
    where $\mathbf{C}(\bm{w})$ denotes the transaction cost function, often proportional or quadratic in the change of portfolio weights. This model demonstrates that incorporating transaction costs ex ante leads to empirically superior and more stable portfolios, especially under model uncertainty.
Each of these models contributes essential insights: the classical mean-variance framework provides the theoretical foundation, the penalized model addresses high-dimensional estimation risk, and the cost-aware model introduces practical considerations arising in real-world trading environments. Our proposed approach, cost-aware portfolio estimator (CAPE), builds upon these developments by integrating both regularization and dynamic transaction cost modeling in a unified framework, tailored for large-scale portfolio optimization.
@@ 2.2 Cost-aware Portfolio Estimator
We consider an investor who aims to allocate wealth across $p$ risky assets over $m$ periods, assuming that asset returns are stationary within each period $t=1,\ldots,m$. We denote the expected asset returns and the corresponding population covariance matrix in the $t$-th period by $\bm{\mu}_{t}$ and $\bm{\Sigma}_{t}$, respectively. Suppose an investor enters the market at day 1 with an initial wealth, allocates the wealth (constructing the first portfolio) at day $n+1$, where $n\in\mathcal{N}^{+}$ is a fixed number of trading days between two consecutive rebalances, and reallocates the wealth  [^1] For simplicity but without loss of generality, we assume the investor rebalances the portfolio at same frequency, e.g., annually as in Section 4. at the beginning of each of the following m1m-1 consecutive periods, where $m$ is the total number of portfolio formation decisions. At every decision day $d$, $d=n+1,2n+1,\ldots,mn+1$, we use the historical returns $\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}]^{\top}\in\mathbb{R}^{n\times p}$, that contains observed returns of asset $j$ on its $j$-th column to estimate assets return and risk. Figure 1 illustrates the decision process.
@@
and the weights are taken at the beginning of period $t$. In real-world applications, investors tend to allocate wealth across a limited number of assets due to transaction costs, regulatory constraints, or interpretability concerns. To incorporate this preference, we consider the transaction cost under the Markowitz mean-variance framework (1.1). And the optimal cost-aware portfolio at first portfolio construction stage ($t=1$) is defined as
$\bm{w}^{*}_{1}:=\mathop{\arg\min}_{\bm{w}\in\mathbb{R}^{p}}\bm{w}^{\top}\bm{\Sigma}_{1}\bm{w}-\gamma\bm{w}^{\top}\bm{\mu}_{1}+\mathbf{C}_{1}(\bm{w})\quad\text{s.t.}\ \bm{w}^{\top}\textbf{1}=1,\ \|\bm{w}\|_{0}\leq s_{0},$ (2.4)
where $\mathbf{C}_{1}(\bm{w})$ is the transaction cost at first portfolio construction stage stage ($t=1$), depending on the trading activity (changes in portfolio weights) and $s_{0}<\infty$ denotes the maximum number of assets allowed in the portfolio. Similarly, the optimal cost-aware portfolio at reallocation stage ($t\geq 2$) is defined as
$\bm{\delta}^{*}_{t}:=\mathop{\arg\min}_{\bm{\delta}\in\mathbb{R}^{p}}\bm{\delta}^{\top}\bm{\Sigma}_{t}\bm{\delta}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}\bm{\mu}_{t}+\mathbf{C}_{t}(\bm{\delta})\quad\text{s.t.}\ \bm{\delta}^{\top}\textbf{1}=0,\ \|\bm{\delta}\|_{0}\leq s_{0},$ (2.5)
where the weight difference vector $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, $\mathbf{C}_{t}(\bm{w})$ is the transaction cost at reallocation stage ($t\geq 2$), and $\bm{w}^{+}_{t-1}$ means the pre-rebalancing weight as the value of portfolio changes over the holding period, and $\bm{w}^{+}_{t-1}=(f_{n}\circ f_{n-1}\circ\dots\circ f_{1})(\bm{w}_{t-1})$, where $f_{i}(\bm{w}):=\frac{\bm{w}\odot(\mathbf{1}+\bm{R}_{(t-2)n+i})}{1+\bm{w}^{\top}\bm{R}_{(t-2)n+i}}$. This method is thus called the cost-aware portfolio estimator (CAPE) in high dimensions. This paper proposes a new estimator targeting $\bm{w}^{*}_{1}$ and $\bm{\delta}^{*}_{t}$ under the high-dimensional model setting.